OpenStudy (anonymous):

In an electric circuit, the reciprocal of the total capacitance of two capacitors in a series in the sum of reciprocals of the capacitances. Find the total capacitence of two capaciatances of 0.0050 F and 0.0040 F connected in series. The total capacitance of the two capacitances is??

OpenStudy (anonymous):

\[\frac{ 1 }{ C_{eq} }=\frac{ 1 }{ C_1 }+\frac{ 1 }{ C_2 }\]

OpenStudy (anonymous):

\(C_1\) and \(C_2\) are the given capacitances and \(C_{eq}\) is what you're solving for

OpenStudy (anonymous):

So can you help me get my solution to get the answer

OpenStudy (anonymous):

Do i need to divide or multipy?

OpenStudy (anonymous):

you need to add the two fractions, then the answer will be the reciprocal

OpenStudy (anonymous):

So I get .009 when I add 0.0050 and 0.0040 together

OpenStudy (anonymous):

put .005 in for C_1 and put .004 in for C2 so you should have \[\frac{ 1 }{ .005 }+\frac{ 1 }{ .004 }\] add this, then take the reciprocal

OpenStudy (anonymous):

\[\frac{ 1 }{ 450 }\] ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So now what to get my answer??

OpenStudy (anonymous):

1/450 is the answer